OpenStudy (anonymous):

solve the formula for r A=pi*r^2

OpenStudy (anonymous):

@satellite73 any chance you could help me out?

OpenStudy (zzr0ck3r):

+-sqrt(a/(pi)) = r

OpenStudy (cwrw238):

divide both sides by pi then take the square root of both sides

OpenStudy (anonymous):

\[\pi*r*r=A\] \[r*r=A/pi\]

OpenStudy (anonymous):

To get r, square root both sides.

OpenStudy (zzr0ck3r):

dont forget the +-

OpenStudy (anonymous):

\[\sqrt{r*r}=r= \pm \sqrt{A/\pi}\]

jimthompson5910 (jim_thompson5910):

zzr0ck3r, because the radius is a length, it doesn't make sense to have a negative radius so you don't have to worry about the plus/minus

OpenStudy (zzr0ck3r):

we are not told this is for a circle, we are given an equation and thats it. could be for something else.

OpenStudy (cwrw238):

@henpen radius cannot be negative

OpenStudy (zzr0ck3r):

says nothig of radius, just gives formula....

OpenStudy (zzr0ck3r):

its good practice.

OpenStudy (anonymous):

Why 'plus or minus': because when you square both sides back again to get r*r=A/pi, if there was a minus sign in front of the (sqrtr*r), then it would square into a plus.

OpenStudy (zzr0ck3r):

because (-2)^2 = 2^2 = 4

OpenStudy (anonymous):

so final answer is.. ?

OpenStudy (anonymous):

i think i got what you are saying so far

OpenStudy (zzr0ck3r):

in this context the final answer is -+sqrt(A/pi) if we were talking about a circle it would not make sense to talk about a negative radius so it would be sqrt(A/pi)

OpenStudy (zzr0ck3r):

but you said nothing about a circle or a radius...

OpenStudy (anonymous):

okay thanks

OpenStudy (cwrw238):

the formula is for the area of a circle and nothing else - the r can only be positive

OpenStudy (zzr0ck3r):

so is this formula only for a circle b = c*r^2??? we are given variables. we are not told they stand for anything else..... dont assume is my point, as this is math. To say a formula only works for one thing is just silly.

OpenStudy (zzr0ck3r):

the r can only be positive once we aply it to some real world problem. but of course unless you know something I dont. A formula is a formula....not set in stone for what it is used for.

OpenStudy (zzr0ck3r):

the r can only be positive once we apply it to some real world problem involving circles....*